OpenStudy (happinessbreaksbones):

help again D: lol y/(y + 3) = 1/(y + 3)

OpenStudy (happinessbreaksbones):

i'm not good at fractions

jimthompson5910 (jim_thompson5910):

The denominators are the same So the numerators must be equal

OpenStudy (happinessbreaksbones):

okay

jimthompson5910 (jim_thompson5910):

Or you can cross multiply y/(y + 3) = 1/(y + 3) y(y + 3) = 1(y + 3) y^2 + 3y = y + 3 y^2 + 3y - y - 3 = 0 y^2 +2y - 3 = 0 (y+3)(y-1) = 0 y+3=0 or y-1=0 y=-3 or y=1 So the *possible* solutions are y = -3 or y = 1, but when y = -3, we get a division by zero error. So we throw out y = -3 Therefore the only solution is y = 1

OpenStudy (happinessbreaksbones):

oh I forgot all about cross multiplying hah thank you :)

jimthompson5910 (jim_thompson5910):

This backs up what I said at the top "The denominators are the same So the numerators must be equal" So if the denominators are the same (which they are -- they're both y+3), then the numerators y and 1 must be the same aka y = 1

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (happinessbreaksbones):

that makes a lot of sense :) good job

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

despite the fact that you can use this trick, i'd still go with cross multiplication since that works in a lot more cases

OpenStudy (happinessbreaksbones):

can you help me on one more?

OpenStudy (happinessbreaksbones):

(x^2 + 2x)/x * (3x - 6)/(x^2 - 4)

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

x^2 + 2x factors to x(x+2) when you factor out the GCF x 3x-6 factors to 3(x-2) when you factor out the GCF 3 and x^2-4 factors to (x-2)(x+2) by the difference of squares rule ------------------------------------------------------- So we then get... \[\Large \frac{x^2+2x}{x} * \frac{3x-6}{x^2-4}\] \[\Large \frac{x(x+2)}{x} * \frac{3x-6}{x^2-4}\] \[\Large \frac{x(x+2)}{x} * \frac{3(x-2)}{x^2-4}\] \[\Large \frac{x(x+2)}{x} * \frac{3(x-2)}{(x-2)(x+2) }\] which combines to \[\Large \frac{3x(x+2)(x-2)}{x(x-2)(x+2)} \] See anything cancelling?

OpenStudy (happinessbreaksbones):

hmm possibly everything but the 3 on the end?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (happinessbreaksbones):

awesome :)

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^2+2x}{x} * \frac{3x-6}{x^2-4}\] \[\Large \frac{x(x+2)}{x} * \frac{3x-6}{x^2-4}\] \[\Large \frac{x(x+2)}{x} * \frac{3(x-2)}{x^2-4}\] \[\Large \frac{x(x+2)}{x} * \frac{3(x-2)}{(x-2)(x+2) }\] \[\Large \frac{3x(x+2)(x-2)}{x(x-2)(x+2)} \] \[\Large \frac{3x(x+2)(x-2)}{x(x-2)(x+2)} \] \[\Large \frac{3x\cancel{(x+2)}(x-2)}{x(x-2)\cancel{(x+2)}} \] \[\Large \frac{3x(x-2)}{x(x-2)} \] \[\Large \frac{3x\cancel{(x-2)}}{x\cancel{(x-2)}} \] \[\Large \frac{3x}{x} \] \[\Large \frac{3\cancel{x}}{\cancel{x}} \] \[\Large \frac{3}{1} \] \[\Large 3 \] So \[\Large \frac{x^2+2x}{x} * \frac{3x-6}{x^2-4} = 3\] for all allowed values of x.

jimthompson5910 (jim_thompson5910):

oops have one duplicated step, but oh well...lol

OpenStudy (happinessbreaksbones):

lol not a problem :) thanks for all the help you really are awesome

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

thanks for the compliment

OpenStudy (happinessbreaksbones):

super math dude

jimthompson5910 (jim_thompson5910):

lol to the rescue...

OpenStudy (anonymous):

hint: solve by cross multiply

OpenStudy (theviper):

y=1 {Denominators are same}